# A Cosmic Miracle: A Remarkably Luminous Galaxy at Confirmed with JWST

### v1

@@ Appendix A Tests for Systematics
We perform two tests to validate different aspects of the NIRSpec data reduction. First, we compare the NIRSpec fluxes to the NIRCam photometry of the source. We synthesize photometry by convolving the observed spectrum with filter curves. The results are shown in Fig. 11. In all filters, spanning 0.9-5m, the fluxes are in excellent agreement within errors. We emphasize that at no point in the NIRSpec data reduction are the NIRCam fluxes incorporated (to e.g., rescale the data). This consistency is crucial to e.g., robustly estimate the slope or fit the damping wing.
In the second test we consider an alternate form of background subtraction instead of our fiducial âlocalâ subtraction that makes use of the nod offset positions. In the âglobalâ background estimation we make use of the empty sky slitlets on the mask that the source was observed in and then fit these data using a sky template with components representing zodiacal dust and a modified Solar spectrum (see Appendix A.2 from de Graaff et al. 2024c for details). The resulting spectrum is shown in Fig. 12. We note that this form of background subtraction comes with its own set of systematics (for e.g., the sky model template may not perfectly track the observed sky). Nonetheless, it provides a handle on the robustness of the features that are the focus of our work. We see that the prominent features from the fiducial spectrum are present here as well, with the various reported lines appearing at and the Ly break consistent with the same redshift.
@@ Appendix B Ruled Out Low-Redshift Solutions
The galaxy from de Graaff et al. (2024a) is used to illustrate the quiescent case by shifting and resampling the spectrum whereas the best-fit Prospector model that was fit to the NIRCam photometry is shown for the scenario. Both these solutions are firmly ruled out by the sharpness of the observed break in the prism spectrum and dearth of flux at shorter wavelengths clearly seen in the 2D spectrum. Furthermore, no lines corresponding to appear in the spectrum â quiescent galaxies at this epoch often show H in emission due to AGN activity (e.g., Carnall et al., 2024).
The only remaining possible type contaminant from the photometry would have been a dusty, strong emission line source whose emission lines could mimick a flat continuum, as seen in the so-called âSchroedingerâ galaxy (e.g., Arrabal Haro et al., 2023; Naidu et al., 2022a). However, this is obviously not the case for MoM-z14, and all possible low- solutions are ruled out.

### v2

@@ Appendix A Tests for Systematics
We perform two tests to validate different aspects of the NIRSpec data reduction. First, we compare the NIRSpec fluxes to the NIRCam photometry of the source. We synthesize photometry by convolving the observed spectrum with filter curves. The results are shown in Fig. 12. We emphasize that at no point in the NIRSpec data reduction are the NIRCam fluxes incorporated (to e.g., rescale the data). In all filters, spanning 0.9-5m, the fluxes are in excellent agreement () within errors. The PRISM/NIRCam ratio of fluxes is F200W: ; F277W: ; F356W: ; F410M: ; F444W: . We note that in all but two bands (F277W, F356W), the NIRCam photometry itself is relatively uncertain (). In the two bands (F277W, F356W) where the source is well-detected (), the offset is modest (). Fortunately, these are also the bands whose wavelength coverage spans almost all the key features of interest (e.g., emission lines, ) and so we expect there to be impact on the derived physical parameters. This overall consistency we find is crucial to e.g., robustly estimate the slope or fit the damping wing.
In the second test, we consider an alternate form of background subtraction instead of our fiducial “local” subtraction that makes use of the nod offset positions. In the “global” background estimation we make use of the empty sky slitlets on the mask that the source was observed in and then fit these data using a sky template with components representing zodiacal dust and a modified Solar spectrum (see Appendix A.2 from de Graaff et al. 2024a for details). The resulting spectrum is shown in Fig. 13. We note that this form of background subtraction comes with its own set of systematics (for e.g., the sky model template may not perfectly track the observed sky). Nonetheless, it provides a handle on the robustness of the features that are the focus of our work. We see that the prominent features from the fiducial spectrum are present here as well, with the various reported lines appearing at and the Ly break consistent with the same redshift.
@@ Appendix B Ruled Out Low-Redshift Solutions
The galaxy from de Graaff et al. (2024c) is used to illustrate the quiescent case by shifting and resampling the spectrum whereas the best-fit Prospector model that was fit to the NIRCam photometry is shown for the scenario. Both these solutions are firmly ruled out by the sharpness of the observed break in the prism spectrum and dearth of flux at shorter wavelengths clearly seen in the 2D spectrum. Furthermore, no lines corresponding to appear in the spectrum – quiescent galaxies at this epoch often show H in emission due to AGN activity (e.g., Carnall et al., 2024).
So-called “Black Hole Stars” (BH*s; e.g., Naidu et al. 2025; de Graaff et al. 2025) that have been proposed to power Little Red Dots (e.g., Matthee et al., 2024) display extremely strong, sharp Balmer breaks far exceeding the maximum break strength of stellar populations thereby possibly mimicking Lyman breaks. However, these sources show a red continuum and broad Balmer lines that are ruled out by the data.
The only remaining possible type contaminant from the photometry would have been a dusty, strong emission line source whose emission lines could mimick a flat continuum, as seen in the so-called ‘Schroedinger’ galaxy (e.g., Arrabal Haro et al., 2023; Naidu et al., 2022b). However, this is obviously not the case for MoM-z14, and all possible low- solutions are ruled out.